OpenStudy (anonymous):

integration of: dx/(sin^4 x + cos^4 x)

sam (.sam.):

\[\int\limits \left(\cos ^4 x+\sin ^4 x\right) \, dx\]

OpenStudy (anonymous):

no. it is in the denominator

sam (.sam.):

\[\int\limits \frac{1}{\cos ^4 x+\sin ^4 x} \, dx\]

OpenStudy (anonymous):

\[\int\limits_{}^{} dx / ( \sin ^{4} + \cos ^{4} ) \]

sam (.sam.):

try converting it to able to use u-sub... \[\int\limits\limits \frac{1}{\cos ^4 x+\sin ^4 x} \, dx \\ \\ \Large \int\limits\limits \frac{\frac{1}{\cos^4x}}{(\cos ^4 x+\sin ^4 x )\frac{1}{\cos^4x}} \, dx\] \[\int\limits \frac{\sec ^4 x}{\tan ^4 x+1} \, dx\] \[\int\limits \frac{\left(1+\tan ^2 x\right) \left(\sec ^2 x\right)}{1+\tan ^4 x} \, dx \] then let u=tanx du=sec^2x \[\int\limits \frac{u^2+1}{u^4+1} \, du\]

OpenStudy (anonymous):

It could also be useful to consider \[ \sin^4(x) = (1-\cos^2(x))^2 \]

OpenStudy (anonymous):

thanks a lot. got it.